# MCTS-RAG: Enhance Retrieval-Augmented Generation with Monte Carlo Tree Search

### v1

@@ Abstract
MCTS-RAG: Enhance Retrieval-Augmented Generation with
@@ 1 Introduction
On the other hand, RAG has been widely used to solve knowledge-intensive tasks Lewis et al. (2020); Karpukhin et al. (2020); Izacard and Grave (2021), but its effectiveness with small-scale LMs remains limited. small-scale LMs struggle with query formulation and retrieved content comprehension, often generating vague queries and misinterpreting key details Fan et al. (2025). Moreover, existing RAG systems do not dynamically adjust their retrieval strategies based on changing informational or reasoning requirements, which results in unnecessary or repetitive retrieval steps Li et al. (2024); Gao et al. (2024). For example, when answering a multi-hop question like “Which novel inspired the movie that won Best Picture in 1994?”, a standard retrieval system might retrieve documents about Forrest Gump (i.e., Best Picture winner in 1994), but fail to recognize the need for additional reasoning or retrieval steps to establish the connection between Forrest Gump and the novel written by Winston Groom. This limitation arises because small-scale language models often lack the ability to refine queries iteratively and integrate retrieved information into a coherent reasoning process.
@@
To address the aforementioned limitations, we propose MCTS-RAG, a novel framework that integrates MCTS’s reasoning and search capabilities with adaptive retrieval mechanisms. At a high level, MCTS-RAG operates by iteratively refining both retrieval and reasoning through a search-based process. Given a query, it explores multiple reasoning paths, dynamically incorporating retrieval actions at key decision points. Retrieved knowledge is then used to evaluate intermediate states, and beneficial retrieval pathways are reinforced through backpropagation. This structured search mechanism ensures that the model efficiently acquires and utilizes relevant information for more accurate reasoning. In contrast, by integrating retrieval with search-based reasoning, MCTS-RAG is able to systematically explore relevant knowledge and reason over it to obtain the correct answer.
@@
MCTS-RAG has the following key features: Improved reasoning accuracy: New retrieval actions enable SLMs to acquire external knowledge and enhance the quality of question answering (§3.2). Optimized query formulation: The refinement process ensures that each query focuses on specific information needs, improving the effectiveness of retrieval query generation (§3.3). Enhanced retrieval quality: Reflecting on and summarizing retrieved information helps reduce semantic discrepancies and ensures alignment with the core problem (§3.3). MCTS-RAG demonstrates superior performance on various knowledge-intensive benchmarks, including ComplexWebQA (CMQA) Talmor and Berant (2018), GPQA Rein et al. (2024), and FoolMeTwice (FMT) Eisenschlos et al. (2021a). Specifically, it achieves over 20% improvement with Llama 3.1-8B and  6% with Qwen2.5-7B on CWQA, roughly 15% and 10% gains on GPQA, and over 10% (Llama) and 4% (Qwen) on FMT, while outperforming other baselines like Standard RAG, ReAct Yao et al. (2023b), Self-Ask Press et al. (2023), Search-O1 Li et al. (2025), and rStar Qi et al. (2024) by effectively retrieving and integrating evidence through refined multi-step reasoning that minimizes hallucinations.
@@ 2 Related Work
The RAG system enhances LLMs in knowledge-intensive tasks by incorporating external information. Query optimization techniques, including expansion and transformation, improve retrieval quality Ma et al. ; Jagerman et al. (2023). Iterative retrieval methods, such as IRCoT Trivedi et al. (2023) and ITER-RETGEN Shao et al. , refine retrieval and generation. LLM-driven retrieval strategies, such as WebGPT Nakano et al. (2021) and Toolformer Schick et al. (2023), have demonstrated notable improvements in efficiency by leveraging large language models to interact with external tools or search engines, thus streamlining the process of gathering relevant data. Meanwhile, self-reflection mechanisms in systems like Self-RAG Asai et al. ; Islam et al. (2024) and Auto-RAG Yu et al. (2024) further enhance retrieval relevance by employing iterative introspection to refine intermediate outputs. To address this, reasoning-intensive retrieval methods have emerged. For example, BRIGHT Su et al. (2024) introduces complex, reasoning-driven queries that challenge traditional retrieval approaches, while Rank1 Weller et al. (2025) leverages advanced inference-time reranking to identify nuanced relationships missed by standard methods. Despite these advancements, however, these methods often overlook alternative solutions due to their linear reasoning approach and the limited capabilities of small-scale LMs.
@@ 3.1 Preliminaries
However, the original rStar framework is limited by its inability to dynamically acquire external knowledge, restricting its performance in knowledge-intensive queries. To address the inherent limitations of rStar, we propose an integrated reasoning framework that combines the iterative reasoning capabilities of rStar with RAG. At a high level, our approach builds on the iterative generative-discriminative structure of rStar and introduces additional operations specifically designed to facilitate dynamic external knowledge retrieval. This enables the language model to seamlessly integrate relevant external information into its reasoning process, significantly improving factual accuracy and decision robustness. The following subsections detail the proposed MCTS-RAG framework.
@@ 6. A6:
Each action is designed to address specific aspects of the reasoning-retrieval interplay, ensuring that the model can adapt its strategy dynamically as it navigates through the problem space. To further enhance exploration, we employ Upper Confidence Bound for Trees (UCT) Kocsis and Szepesvári (2006) in our MCTS framework—a crucial method that balances exploitation and exploration. The UCT formula is:
$\text{UCT}(s,a)=\bar{Q}(s,a)+C\cdot\sqrt{\frac{\ln N(s)}{N(s,a)}},$
@@ 3.4 Determing Final Answer
ensuring that the most frequent and consistent reasoning trajectory is chosen.
4 Experiments
4.1 Experimental Setup
We evaluate Qwen2.5-7B and Llama 3.1-8B on three complex reasoning tasks: ComplexWebQA (CWQA) Talmor and Berant (2018), which requires multi-step reasoning over web-based queries; Graduate-Level Google-Proof QA (GPQA) Rein et al. (2023), which tests knowledge-intensive science question answering; and FoolMeTwice (FMT) Eisenschlos et al. (2021b), a challenging fact-checking benchmark that assesses the model’s ability to verify factual claims.
@@ 4.1 Experimental Setup
CoT 34.65 35.00 57.25 27.72 28.71 56.50
GPT-4o 54.45 52.98 55.44 54.45 52.98 55.44
Qwen2.5-72B 44.55 40.59 58.41 44.55 40.59 58.41
rStar 55.45 32.32 55.94 37.62 28.71 56.42
@@
GPT-4o 59.40 54.90 61.38 59.40 54.90 61.38
Qwen2.5-72B 48.51 43.13 59.40 48.51 43.13 59.40
ReAct 45.54 41.58 62.37 47.52 34.31 55.44
Self-Ask 44.55 42.57 60.91 44.55 57.84 58.41
Search-O1 49.50 54.45 64.35 44.55 58.82 62.87
MCTS-RAG 61.38 64.64 68.28 67.32 74.25 74.25
Table 1: Answer accuracy of MCTS-RAG and other methods (both with and without retrieval modules).
Settings CWQA GPQA FMT
@@
12 rollout 68.65 75.15 69.35
16 rollout 71.20 84.31 74.14
Table 2: Answer accuracy of Qwen2.5-7B-Instruct under different retrieval and rollout settings.
Baselines.
We evaluate the following baseline methods for comparison: Chain-of-Thought (CoT) Wei et al. (2022) prompting encourages the model to generate explicit step-by-step reasoning to improve complex problem-solving. Standard RAG Ma et al. (2023) performs a single-pass retrieval to augment the model’s responses but lacks iterative refinement. ReAct Yao et al. (2023b) alternates between reasoning and retrieval, allowing the model to dynamically refine its understanding based on external evidence. Self-Ask with Search (Self-Ask) Press et al. (2023) with Search decomposes complex queries into subquestions, retrieves relevant external information, and synthesizes the answers to enhance multi-step reasoning. Search-O1 Li et al. (2025) executes a single retrieval step before generating an answer, limiting its ability to iteratively verify information. Finally, the original rStar Qi et al. (2024) algorithm employs a structured iterative reasoning process but does not fully leverage dynamic retrieval or decomposition.
Implementation Details.
We compare MCTS-RAG with other baselines using the same language models, Qwen2.5-7B and LLaMA 3.1-8B, to ensure a fair comparison of reasoning capabilities. To maintain consistency across methods, we use the same retrieval corpus and retriever, thus avoiding discrepancies caused by varying information access. Specifically, we rely on the Bing Search Engine and LangChain for retrieval, as Bing provides extensive and up-to-date web information while LangChain supports robust retrieval-augmented generation workflows. For different datasets, CWQA is obtained from its web snippets, GPQA uses a corpus sourced from Wikipedia and Bing, and FMT draws on its own related documents. This setup ensures that variations in performance stem from differences in reasoning mechanisms. To facilitate structured reasoning, we configure our setup with a rollout of 4, allowing multiple steps of reasoning expansion. Each query can be decomposed into at most two subquestions, ensuring a controlled breakdown of complex queries. We set the maximum reasoning depth to 5, enabling deep but efficient multi-hop reasoning.
4.2 Main Findings
Table 1 compares reasoning methods on CWQA, GPQA, and FMT for Llama 3.1-8B and Qwen2.5-7B. Our approach consistently outperforms baselines, demonstrating strong multi-step reasoning and retrieval capabilities. On CWQA, it achieves over a 20% gain with Llama 3.1-8B and around 6% with Qwen2.5-7B. Similarly, it surpasses competitors on GPQA by roughly 15% and 10%, respectively, benefiting from refined verification strategies. On FMT, it leads by over 10% with Llama 3.1-8B and 4% with Qwen2.5-7B, proving its resilience against misleading distractors. These results highlight our method’s superior generalization and efficiency, especially in fact-checking and science-related tasks. Compared to baselines like Standard RAG, ReAct, Self-Ask, and Search-O1, Our structured multi-step reasoning can retrieve and process evidence more accurately, and on average we improve the performance by about 14% over the baseline under three datasets. Unlike rStar, it enables broader retrieval, extracting critical insights while minimizing hallucinations, achieving an average improvement of 17%. This framework sets a new benchmark for complex reasoning, delivering high accuracy and efficiency in diverse problem-solving scenarios.

### v2

# MCTS-RAG: Enhancing Retrieval-Augmented Generation with Monte Carlo Tree Search

@@ Abstract
MCTS-RAG: Enhancing Retrieval-Augmented Generation with
@@ 1 Introduction
Figure 1: Overview of the research. The top panel illustrates the proposed MCTS-RAG framework, while the bottom panel summarizes three key findings from our experiments and analysis.
@@
On the other hand, RAG has been widely used to solve knowledge-intensive tasks Lewis et al. (2020); Karpukhin et al. (2020); Izacard and Grave (2021); Zhao et al. (2025), but its effectiveness with small-scale LMs remains limited. small-scale LMs struggle with query formulation and retrieved content comprehension, often generating vague queries and misinterpreting key details Fan et al. (2025). Moreover, existing RAG systems do not dynamically adjust their retrieval strategies based on changing informational or reasoning requirements, which results in unnecessary or repetitive retrieval steps Li et al. (2024); Gao et al. (2024). For example, when answering a multi-hop question like “Which novel inspired the movie that won Best Picture in 1994?”, a standard retrieval system might retrieve documents about Forrest Gump (i.e., Best Picture winner in 1994), but fail to recognize the need for additional reasoning or retrieval steps to establish the connection between Forrest Gump and the novel written by Winston Groom. This limitation arises because small-scale LMs often lack the ability to refine queries iteratively and integrate retrieved information into a coherent reasoning process.
@@
To address the aforementioned limitations, we propose MCTS-RAG, a novel framework that integrates MCTS’s reasoning and search capabilities with adaptive retrieval mechanisms. At a high level, MCTS-RAG operates by iteratively refining both retrieval and reasoning through a search-based process. Given a query, it explores multiple reasoning paths, dynamically incorporating retrieval actions at key decision points. Retrieved knowledge is then used to evaluate intermediate states, and beneficial retrieval pathways are reinforced through backpropagation. This structured search mechanism ensures that the model efficiently acquires and utilizes relevant information for more accurate reasoning. To further enhance efficiency, MCTS-RAG employs parallel expansion and retrieval pruning strategies during the search, reducing redundant computation while maintaining search quality. In contrast to prior approaches, by integrating retrieval with search-based reasoning, MCTS-RAG is able to systematically explore relevant knowledge and reason over it to obtain the correct answer.
@@
MCTS-RAG has the following key features: Improved reasoning accuracy: New retrieval actions enable SLMs to acquire external knowledge and enhance the quality of question answering (§3.2). Optimized query formulation: The refinement process ensures that each query focuses on specific information needs, improving the effectiveness of retrieval query generation (§3.3). Enhanced retrieval quality: Reflecting on and summarizing retrieved information helps reduce semantic discrepancies and ensures alignment with the core problem (§3.3). High efficiency: Parallel expansion and retrieval pruning reduce redundant computation during search, greatly improving inference efficiency without compromising performance (§4.5).
MCTS-RAG demonstrates superior performance on various knowledge-intensive benchmarks, including ComplexWebQA (CMQA) Talmor and Berant (2018), GPQA Rein et al. (2024), and FoolMeTwice (FMT) Eisenschlos et al. (2021a). Specifically, it achieves over 20% improvement with Llama 3.1-8B and  6% with Qwen2.5-7B on CWQA, roughly 15% and 10% gains on GPQA, and over 10% (Llama) and 4% (Qwen) on FMT, while outperforming other competitive baselines. Our efficiency analysis demonstrates that MCTS-RAG delivers the best overall trade-off compared to other baseline systems, achieving the highest accuracy with moderate latency and token cost.
@@ 2 Related Work
The RAG system enhances LLMs in knowledge-intensive tasks by incorporating external information. Query optimization techniques, including expansion and transformation, improve retrieval quality Ma et al. ; Jagerman et al. (2023). Iterative retrieval methods, such as IRCoT Trivedi et al. (2023) and ITER-RETGEN Shao et al. (2023), refine retrieval and generation. LLM-driven retrieval strategies, such as WebGPT Nakano et al. (2021) and Toolformer Schick et al. (2023), have demonstrated notable improvements in efficiency by leveraging large language models to interact with external tools or search engines, thus streamlining the process of gathering relevant data. Meanwhile, self-reflection mechanisms in systems like Self-RAG Asai et al. ; Islam et al. (2024) and Auto-RAG Yu et al. (2024) further enhance retrieval relevance by employing iterative introspection to refine intermediate outputs. More recently, ReARTeR Sun et al. (2025) improves retrieval by reasoning over multiple answer candidates with mutual evaluation. However, its sequential processing and full candidate scoring lead to high latency. RAG-Star Jiang et al. (2024a) uses a fixed search tree for token-level retrieval control, but lacks dynamic pruning and does not support concurrent reasoning paths. In contrast, MCTS-RAG supports parallel expansion of diverse reasoning strategies and incorporates lightweight retrieval pruning to skip unnecessary external calls. This leads to better efficiency without sacrificing answer quality.
@@ 3.1 Preliminaries
[…]
@@ 6. A6:
Algorithm 1 illustrates the procedure for updating each action’s reward. To further enhance exploration, we employ Upper Confidence Bound for Trees (UCT) Kocsis and Szepesvári (2006) in our MCTS framework—a crucial method that balances exploitation and exploration. The UCT formula is:
$\text{UCT}(s,a)=\bar{Q}(s,a)+C\cdot\sqrt{\frac{\ln N(s)}{N(s,a)}}$
@@
Algorithm 1 $R(s,a)$ Computation and Update
1:State $s$, action $a$, completions $\{o_{1},\dots,o_{K}\}$, log-likelihoods $\{\ell_{1},\dots,\ell_{K}\}$
2:Representative answer $o^{*}$, confidence $\mathrm{Conf}(o^{*})$, reward $R(s,a)$
3:Initialize empty clusters $\mathcal{C}\leftarrow\emptyset$
4:for $j=1$ to $K$ do
5:  $o_{j}\leftarrow$ $j$-th completion
6:  matched $\leftarrow$ False
7:  for each representative $r$ in $\mathcal{C}$ do
8:   if $\texttt{Equiv}(o_{j},r)$ then
9:     Add $o_{j}$ to cluster $\mathcal{C}_{r}$
10:     matched $\leftarrow$ True
11:     break
12:   end if
13:  end for
14:  if not matched then
15:   Create new cluster $\mathcal{C}_{o_{j}}\leftarrow\{o_{j}\}$
16:  end if
17:end for
18:Let $\mathcal{O}^{*}\leftarrow$ majority cluster with size $n^{*}$
19:Select representative $o^{*}\in\mathcal{O}^{*}$
[…]
@@ 3.4 Determing Final Answer
ensuring that the most frequent and consistent reasoning trajectory is chosen. Essentially, our approach operates as a reward‐accumulation voting scheme: by normalizing and aggregating the trajectory rewards of semantically consistent candidate clusters, it guarantees that the selected answer is both the most coherent and the highest‐scoring.
4 Experiment Setup
4.1 Evaluation Benchmark
We evaluate MCTS-RAG and other competitive baseline systems on three complex reasoning tasks: (1) ComplexWebQA (CWQA) Talmor and Berant (2018), which requires multi-step reasoning over web-based queries; (2) GPQA Rein et al. (2023), which tests knowledge-intensive science question answering; and (3) FoolMeTwice (FMT) Eisenschlos et al. (2021b), a challenging fact-checking benchmark that assesses the model’s ability to verify factual claims.
4.2 Baseline Systems
Beyond CoT prompting, rStar, and Standard RAG, we also compare MCTS-RAG with several recent RAG variants, including: ReAct Yao et al. (2023b) alternates between reasoning and retrieval, allowing the model to dynamically refine its understanding based on external evidence. Self-Ask with Search (Self-Ask) Press et al. (2023) with Search decomposes complex queries into subquestions, retrieves relevant external information, and synthesizes the answers to enhance multi-step reasoning. Search-O1 Li et al. (2025) executes a single retrieval step before generating an answer, limiting its ability to iteratively verify information. Self-RAG Asai et al. (2024) extends standard RAG by generating and issuing its own retrieval queries at each decoding step, enabling deeper multi-hop evidence gathering. FLARE Jiang et al. (2023) employs an active learning strategy to select the most informative passages for retrieval, improving answer relevance with fewer retrieval calls. Iter-RetGen Shao et al. (2023) alternates retrieval and generation phases in multiple passes, refining responses through iterative editing. AutoRAG Kim et al. (2024) automates retrieval pipeline configuration and model selection to optimize end-to-end retrieval-augmented generation performance. TC-RAG Jiang et al. (2024b) integrates Turing-complete control flows into RAG, supporting complex multi-step reasoning via case-based retrieval loops. IterDRAG Yue et al. introduces dynamic retrieval–generation loops that adapt retrieval strategies based on intermediate model outputs. DeepRAG Guan et al. (2025) interleaves deep reasoning modules between retrieval steps to enhance multi-hop inference capabilities. ReARTeR Sun et al. (2025) uses a reward model and explanations to improve step-by-step reasoning with MCTS. RAG-Star Jiang et al. (2024a) combines MCTS planning with external retrieval to guide and verify multi-step reasoning. To ensure fair comparisons, we use the same base LLMs—Qwen2.5-7B and Llama 3.1-8B—for all the evaluated systems.
4.3 Implementation Details
RAG Setup.
To maintain consistency across methods, we use a shared retrieval corpus and identical retriever configurations. We employ the Bing Search Engine and LangChain for retrieval, with Bing offering extensive and up-to-date web information and LangChain providing modular support for retrieval-augmented generation workflows. For the CWQA dataset, we collect approximately 100K web snippets retrieved via Bing using question templates; these snippets are dynamically retrieved at inference time based on the input question. For GPQA, we construct a static corpus comprising 80K passages sampled from Wikipedia and 60K web documents retrieved from Bing, totaling 140K documents. The retriever encodes questions and selects top-ranked documents from this hybrid source. For FMT, a domain-specific dataset, we use its original associated documents (about 30K passages) provided as part of the benchmark and treat them as the retrieval pool. In all cases, top-10 retrieved documents are fed into the reasoning module for answer generation and verification. This setup ensures that variations in performance stem from differences in reasoning mechanisms.
MCTS-RAG Setup.
To facilitate structured reasoning, we configure our setup with a rollout of 4, allowing multiple steps of reasoning expansion. Each query can be decomposed into at most two subquestions, ensuring a controlled breakdown of complex queries. We set the maximum reasoning depth to 5, enabling deep but efficient multi-hop reasoning. Moreover, in order to reduce latency during tree expansion, we implement concurrent action evaluations: different reasoning actions at a given search node are expanded in parallel. This design leverages asynchronous computation to significantly speed up MCTS traversal without sacrificing the fidelity of action-value estimation.
@@ 4.3 Implementation Details
CoT 34.6 35.0 57.3 27.7 28.7 56.5
GPT-4o 54.4 53.0 55.4 54.5 53.0 55.4
Qwen2.5-72B 44.5 40.6 58.4 44.6 40.6 58.4
rStar 55.4 32.3 55.9 37.6 28.7 56.4
Standard RAG 44.2 40.6 58.4 35.6 31.7 51.5
GPT-4o 59.4 54.9 61.4 59.4 54.9 61.4
Qwen2.5-72B 48.5 43.1 59.4 48.8 46.2 59.8
ReAct 45.5 41.6 62.4 47.5 49.3 55.4
Self-Ask 44.6 42.6 60.9 44.6 52.8 58.4
Self-RAG 46.2 43.1 61.1 47.1 53.9 60.2
FLARE 50.1 45.3 62.6 50.1 56.6 62.1
ReARTeR 51.8 46.4 63.3 51.4 57.1 64.3
Iter-RetGen 52.2 47.5 63.1 52.3 57.6 63.2
Search-O1 49.5 54.5 64.4 54.6 58.8 65.9
AutoRAG 57.2 55.1 66.1 57.2 59.3 66.9
TC-RAG 57.7 55.5 66.7 59.6 62.3 68.7
RAG-Star 58.1 57.9 67.4 60.1 66.3 69.8
IterDRAG 59.3 58.2 67.1 61.1 67.2 71.1
DeepRAG 60.9 61.3 66.9 62.3 69.8 72.9
MCTS-RAG 61.4 64.6 68.3 67.3 71.3 73.8
Table 1: Answer accuracy of MCTS-RAG and other methods (with and without retrieval modules).
4.4 Main Findings
Table 1 compares reasoning methods on CWQA, GPQA, and FMT for Llama 3.1-8B and Qwen2.5-7B. Our approach consistently outperforms baselines, demonstrating strong multi-step reasoning and retrieval capabilities. On CWQA, it achieves over a 20% gain with Llama 3.1-8B and around 6% with Qwen2.5-7B. Similarly, it surpasses competitors on GPQA by roughly 42% and 32%, respectively, benefiting from refined verification strategies. On FMT, it leads by over 17% with Llama 3.1-8B and 12% with Qwen2.5-7B, proving its resilience against misleading distractors. These results highlight our method’s superior generalization and efficiency, especially in fact-checking and science-related tasks. Compared to baselines like Standard RAG, ReAct, Self-Ask, Search-O1, TC-RAG, IterDRAG, and DeepRAG, our structured multi-step reasoning can retrieve and process evidence more accurately, and on average we improve the performance by about 14% over the baseline under three datasets. Compared to rStar, MCTS-RAG enables broader retrieval, extracting critical insights while minimizing hallucinations, achieving an average improvement of 17%.
4.5 Efficiency Analysis
Figure 4: Comparison of different methods on GPQA using Qwen2.5-7B. The horizontal axis shows accuracy, the vertical axis shows relative latency the decoding time relative to the Standard RAG baseline, and marker size and color correspond to average token generated.
As shown in Figure 4, the standard RAG system serves as the baseline with normalized latency (1.0×) but also the lowest accuracy (40.6%). ReAct improves accuracy to 42.6% at the cost of 2.0× latency and slightly lower token consumption (8,884). Self-RAG further increases accuracy to 43.1% with 2.8× latency and 11,000 tokens. Search-O1 balances accuracy (54.5%) and latency (2.8×) with moderate token cost (11,300), making it suitable for latency-sensitive scenarios. IterDRAG (56.1%, 3.3×, 12,500 tokens) and DeepRAG (57.3%, 4.0×, 13,700 tokens) trade additional latency for marginal accuracy gains. rStar performs poorly (32.3%, 5.5×, 19,000 tokens), indicating low efficiency. MCTS-RAG delivers the best overall trade-off, achieving the highest accuracy (64.6%) with moderate latency (2.8×) and token cost (11,892).
@@ 4.5 Efficiency Analysis
This improvement stems from two key optimizations in MCTS-RAG: (1) Parallel Expansion of Reasoning Actions: during each rollout step, candidate subquestions (“actions”) are generated and verified in parallel, rather than sequentially as in rStar. This concurrency amortizes the overhead of verification across multiple branches, yielding higher token throughput per unit time. (2) Dynamic Pruning of Unnecessary Retrievals: Our framework allows the model to autonomously determine whether a retrieval step is needed. If the current context suffices, the model skips retrieval entirely. To improve both accuracy and efficiency, this pruning decision is guided by two lightweight yet effective signals: (i) a retrieval necessity signal, where the model first decides, based on the current prompt and context, whether external retrieval is required, thus avoiding unnecessary queries; and (ii) a candidate similarity signa, where multiple reasoning candidates are compared at each step and branches with abnormally low consistency are pruned, as they typically indicate hallucinations or reasoning failures. This implicit gating mechanism reduces redundant queries without introducing additional pruning modules. Together, these signals enhance reasoning stability, reduce computational cost, and are particularly valuable in resource-constrained environments.
[…]
Settings CWQA GPQA FMT Token Latency
@@
[…]
Table 2: Accuracy of Qwen2.5-7B under various retrieval and rollout settings. Token: avg. generated tokens; Latency: relative decoding time vs standard RAG.